OpenStudy (anonymous):

Which is a solution to the equation y = 2x – 5? A.(3, 11) B.(2, 1) C.(1, –3) D.(5, –5)

Mehek (mehek14):

which one do you think?

OpenStudy (korosh23):

Insert the coordinates inside the original equation, Pick everyone and see if they work, but for you to learn I am going to give you an example. y= 3x +2 (2,2) (3,11) (2) = 3(2) +2 2= 8 No, because they do not equal to each other, so reject the coordinate. Now try this: (11) = 3 (3) +2 11= 9 + 2 11= 11 Correct, this is the solution to the equation y= 3x+2 This was just an example, follow this and then you can solve your question! :) I hope this helped, and if you had any question, feel free to let me know.